Equivariant Operational Chow Rings

of T-Linear Schemes

Richard P. Gonzales^{1}

Received: May 14, 2014 Revised: December 26, 2014

Communicated by Alexander Merkurjev

Abstract. We study T -linear schemes, a class of objects that includes spherical and Schubert varieties. We provide a localization theorem for the equivariant Chow cohomology of these schemes that does not depend on resolution of singularities. Furthermore, we give an explicit presentation of the equivariant Chow cohomology of possibly singular complete spherical varieties admitting a smooth equivariant envelope (e.g., group embeddings).

2010 Mathematics Subject Classification: 14M27, 14L30, 20M32 Keywords and Phrases: spherical varieties, Chow cohomology, Kronecker duality, intersection theory.

1 Introduction and motivation

Let 𝕜 be an algebraically closed field. Let G be a connected reductive linear algebraic group (over 𝕜). Let B be a Borel subgroup of G and T ⊂ B be a maximal torus of G. An algebraic variety X, equipped with an action of G, is spherical if it contains a dense orbit of B. (Usually spherical varieties are assumed to be normal but this condition is not needed here.) Spherical varieties have been extensively studied in the works of Akhiezer, Brion, Knop, Luna, Pauer, Vinberg, Vust and others. For an up-to-date discussion of spherical varieties, as well as a comprehensive bibliography, see [Ti] and the references therein. If X is spherical, then it has a finite number of B-orbits, and thus, also a finite number of G-orbits [Ti]. In particular, T acts on X with a finite number of fixed points. These properties make spherical varieties particularly well suited for applying the methods of Goresky-Kottwitz-MacPherson [GKM], nowadays called GKM theory, in the topological setup, and Brion’s extension of GKM theory [Br2] to the algebraic setting of equivariant Chow groups, as defined by Totaro, Edidin and Graham [EG1]. Through this method, substantial information about the topology and geometry of a spherical variety can be obtained by restricting one’s attention to the induced action of T. Examples of spherical varieties include G×G-equivariant embeddings of G (e.g., toric varieties are spherical) and the regular symmetric varieties of De Concini-Procesi [DP]. The equivariant cohomology and equivariant Chow groups of smooth complete spherical varieties have been studied by Bifet, De Concini and Procesi [BCP], De Concini-Littelmann [LP], Brion [Br2] and Brion-Joshua [BJ2]. In these cases, there is a comparison result relating equivariant cohomology with equivariant Chow groups: for a smooth complete spherical variety, the equivariant cycle map yields an isomorphism from the equivariant Chow group to the equivariant (integral) cohomology (Proposition 2.11). As for the study of the equivariant Chow groups of possibly singular spherical varieties, some progress has been made by Payne [P] and the author [G2]-[G4]. The problem of developing intersection theory on singular varieties comes from the fact that the Chow groups A_{∗}(−) do not admit, in general, a natural ring structure or intersection product. But when singularities are mild, for instance when X is a quotient of a smooth variety Y by a finite group F, then A_{∗}(X) ⊗ ℚ ≃ (A_{∗}(Y) ⊗ ℚ)^{F}, and so A_{∗}(X) ⊗ ℚ inherits the ring structure of A_{∗}(Y) ⊗ ℚ. To simplify notation, if A is a ℤ-module, we shall write hereafter A_{ℚ} for the rational vector space A ⊗ ℚ. In order to study more general singular schemes, Fulton and MacPherson [Fu] introduced the notion of operational Chow groups or Chow cohomology. Similarly, Edidin and Graham defined the equivariant operational Chow groups [EG1], which we briefly recall. (For our conventions on varieties and schemes, see Section 2.1.) Let X be a T -scheme. The i-th T -equivariant operational Chow group of X, denoted A_{T}^{i} (X), is defined as follows. An element c ∈ A_{T}^{i} (X) is a collection of homomorphisms c_{f}^{(m)} : A_{m}^{T}(Y) → A_{m−i}^{T}(Y), written z ↦ f^{∗}c ∩ z, for every T -equivariant map f : Y → X and all integers m (the underlying category is the category of T -schemes). Here A_{∗}^{T} (Y) denotes the equivariant Chow group of Y (Section 2.1). As in the case of ordinary operational Chow groups, these homomorphisms must satisfy three conditions of compatibility: with proper pushforward (resp. flat pull-back, resp. intersection with a Cartier divisor) for T -equivariant maps Y′ → Y → X, with Y′ → Y proper (resp. flat, resp. determined by intersection with a Cartier divisor); see [Fu, Chapter 17] for precise statements. The homomorphism c_{f}^{(m)} determined by an element c ∈ A_{T}^{i} (X) is usually denoted simply by c, with an indication of where it acts. For any X, the ring structure on A_{T}^{∗} (X) := ⊕_{i}A_{T}^{i} (X) is given by composition of such homomorphisms. The ring A_{T}^{∗} (X) is graded, and A_{T}^{i} (X) can be non-zero for any i ≥ 0. The most salient functorial properties of equivariant operational Chow groups are summarized below:

1Supported by the Institut des Hautes Études Scientifiques, TÜBİTAK Project No.

112T233, and DFG Research Grant PE2165/1-1.

Documenta Mathematica 20 (2015) 401–432

(i) Cup products A_{T}^{p} (X) ⊗ A_{T}^{q} (X) → A_{T}^{p+q}(X), a ⊗ b ↦ a ∪ b, making A_{T}^{∗} (X) into a graded associative ring (commutative when resolution of singularities is known).

(ii) Contravariant graded ring maps f^{∗} : A_{T}^{i} (X) → A_{T}^{i} (Y) for arbitrary equivariant morphisms f : Y → X.

(iii) Cap products A_{T}^{i} (X)⊗A_{m}^{T}(X) → A_{m−i}^{T}(X), c⊗z ↦ c∩z, making A_{∗}^{T} (X) into an A_{T}^{∗} (X)-module and satisfying the projection formula.

(iv) If X is a nonsingular n-dimensional T -variety, then the Poincaré duality map from A_{T}^{i} (X) to A_{n−i}^{T}(X), taking c to c∩[X], is an isomorphism, and the ring structure on A_{T}^{∗} (X) is that determined by intersection products of cycles on the mixed spaces X_{T} [EG1, Proposition 4].

(v) Equivariant vector bundles on X have equivariant Chern classes in A_{T}^{∗} (X).

(vi) Localization theorems of Borel-Atiyah-Segal type and GKM theory (with rational coefficients) for possibly singular complete T -varieties in characteristic zero. See [G3] or the Appendix for details.

In [FMSS], Fulton, MacPherson, Sottile and Sturmfels succeed in describing the non-equivariant operational Chow groups of complete spherical varieties. Indeed, they show that the Kronecker duality homomorphism

K : A^{i}(X) ⟶ Hom(A_{i}(X), ℤ), α ↦ (β ↦ deg(β ∩ α))

is an isomorphism for complete spherical varieties. Here deg (−) is the degree homomorphism A_{0}(X) → ℤ. Moreover, they prove that A_{∗}(X) is finitely generated by the classes of B-orbit closures, and with the aid of the map K, they provide a combinatorial description of A^{∗}(X) and the structure constants of the cap and cup products [FMSS]. In addition, if X is nonsingular and complete, they show that the cycle map cl_{X} : A_{∗}(X) → H_{∗}(X) is an isomorphism. Although we stated their results in the case of spherical varieties, these hold more generally for complete schemes with a finite number of orbits of a solvable group. In particular, the conclusions of [FMSS] hold for Schubert varieties. Later on, Totaro [To] extended these results to the broader class of linear schemes, a class first studied in work of Jannsen [Ja]. The results of [FMSS] and [To] are quite marvelous in that they give a presentation of a rather abstract ring, namely A^{∗}(X), in a very combinatorial manner.

In this article, we extend the results of the previous paragraph to the equivariant Chow cohomology of T -linear schemes (Definition 2.3). By Theorem 2.5, spherical varieties are T -linear (this fact does not follow directly from [To] and [R], see the comments before Theorem 2.5). Also, we obtain a localization theorem for the equivariant Chow cohomology of complete T -linear schemes that does not depend on resolution of singularities (Theorem 3.9). Last, and most important, we give a presentation of the rational equivariant Chow cohomology of complete possibly singular spherical varieties admitting an equivariant smooth envelope (Theorem 4.8). The latter vastly increases the applicability of Brion’s techniques [Br2, Section 7] from the smooth to the singular setup. Here is an outline of the paper. Section 2 reviews the necessary background material. Section 3 is the conceptual core of this article. In Subsection 3.1 we obtain the equivariant versions of the results of [FMSS] and [To] that concern us. We start by defining equivariant Kronecker duality schemes. These are complete T -schemes X which satisfy two conditions: (i) A_{∗}^{T} (X) is finitely generated over S = A_{T}^{∗} (pt), and (ii) the equivariant Kronecker duality map K_{T} : A_{T}^{∗} (X) ⟶ Hom_{S}(A_{∗}^{T} (X), S) is an isomorphism of S-modules (Definition 3.1). As an example, we show that complete T -linear schemes satisfy equivariant Kronecker duality (Proposition 3.5). This is deduced from the equivariant Künneth formula (Proposition 3.4). In Subsection 3.2 we prove our second main result, namely, a localization theorem for equivariant Kronecker duality schemes (Theorem 3.9). We conclude Section 3 by showing that projective group embeddings in arbitrary characteristic satisfy equivariant localization (Theorem 3.11). This extends well-known results on torus embeddings [P] to more general compactifications of connected reductive groups. Finally, in Section 4, we apply the machinery just developed to spherical varieties in characteristic zero and prove the most important result of this paper, namely, Theorem 4.8. It asserts that if X is a complete, possibly singular, G-spherical variety, then the image of the injective map i_{T}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (X^{T})_{ℚ} is fully described by congruences involving pairs, triples or quadruples of T -fixed points. Remarkably, this extends [Br2, Theorem 7.3] to the singular setting.

Acknowledgments

Most of the research in this paper was done during my first visit to the Institute des Hautes Études Scientifiques (IHES). I am deeply grateful to IHES for its support, outstanding hospitality and excellent atmosphere. A very special thank you goes to Michel Brion for carefully reading previous drafts of this article, and for offering detailed comments and suggestions on various parts. I would also like to thank Sabancı Üniversitesi, the Scientific and Technological Research Council of Turkey (TÜBİTAK), and the German Research Foundation (DFG) for their support during the final stages of completing this article. I am further grateful to the referee for very helpful comments and suggestions that improved the presentation and scope of the article.

2 Definitions and basic properties

2.1 Conventions and notation

Throughout this paper, we fix an algebraically closed field 𝕜 (of arbitrary characteristic, unless stated otherwise). All schemes and algebraic groups are assumed to be defined over 𝕜. By a scheme we mean a separated scheme of finite type over 𝕜. A variety is a reduced and irreducible scheme. A subvariety is a closed subscheme which is a variety. A point on a scheme will always be a closed point. The additive and multiplicative groups over 𝕜 are denoted by 𝔾_{a} and 𝔾_{m}. We denote by T an algebraic torus. We write Δ for the character group of T, and S for the symmetric algebra over ℤ of the abelian group Δ. We denote by Q the quotient field of S. A scheme X provided with an algebraic action of T is called a T -scheme. For a T -scheme X, we denote by X^{T} the fixed point subscheme and by i_{T} : X^{T} → X the natural inclusion. If H is a closed subgroup of T, we similarly denote by i_{H} : X^{H} → X the inclusion of the fixed point subscheme. When comparing X^{T} and X^{H} we write i_{T,H} : X^{T} → X^{H} for the natural (T -equivariant) inclusion.

A T -scheme X is called locally linearizable (and the T -action is called locally linear) if X is covered by invariant affine open subsets. For instance, T -stable subschemes of normal T -schemes are locally linearizable [Su]. A T -scheme is called T -quasiprojective if it has an ample T -linearized invertible sheaf. This assumption is satisfied, e.g. for T -stable subschemes of normal quasiprojective T -schemes [Su]. Recall that an envelope p : X̃ → X is a proper map such that for any subvariety W ⊂ X there is a subvariety W̃ mapping birationally to W via p [Fu, Definition 18.3]. In the case of T -actions, we say that p : X̃ → X is an equivariant envelope if p is T -equivariant, and if we can take W̃ to be T - invariant for T -invariant W. If there is a dense open set U ⊂ X over which p is an isomorphism, then we say that p : X̃ → X is a birational envelope. By [Su, Theorem 2], if X is a T -scheme, then there exists a T -equivariant birational envelope p : X̃ → X, where X̃ is a T -quasiprojective scheme. Moreover, if char(𝕜) = 0, then we may choose X̃ to be smooth [EG2, Proposition 7.5]. If p : X̃ → X is a T -equivariant envelope, and H ⊂ T is a closed subgroup, then the induced map X̃^{H} → X^{H} is a T -equivariant envelope [EG2, Lemma 7.2].

Let X be a T -scheme of dimension n (not necessarily equidimensional). Let V be a finite dimensional T -module, and let U ⊂ V be an invariant open subset such that a principal bundle quotient U → U/T exists. Then T acts freely on X × U and the quotient scheme X_{T} := (X × U)/T exists. Following Edidin and Graham [EG1], we define the i-th equivariant Chow group A_{i}^{T} (X) by A_{i}^{T} (X) := A_{i+dim U−dim T} (X), if V \ U has codimension more than n − i. Such pairs (V, U) always exist, and the definition is independent of the choice of (V, U), see [EG1]. Finally, A_{∗}^{T} (X) := ⊕_{i}A_{i}^{T} (X). Unlike ordinary Chow groups, A_{i}^{G}(X) can be non-zero for any i ≤ n, including negative i. If X is a T -scheme, and Y ⊂ X is a T -stable closed subscheme, then Y defines a class [Y] in A_{∗}^{T} (X). If X is smooth and equidimensional, then so is X_{T}, and A_{∗}^{T} (X) admits an intersection pairing; in this case, the corresponding ring graded by codimension is isomorphic to the equivariant operational Chow ring A_{T}^{∗} (X) [EG1, Proposition 4]. The equivariant Chow ring of a point A_{T}^{∗} (pt) identifies to S, and A_{∗}^{T} (X) is a S-module, where Δ acts on A_{∗}^{T} (X) by homogeneous maps of degree −1. This module structure is induced by pullback through the flat map p_{X,T} : X_{T} → U/T. Restriction to a fiber of p_{X,T} gives a canonical map A_{∗}^{T} (X) → A_{∗}(X), and this map is surjective (Theorem 2.6). If X is complete, we denote by p_{X,T ∗}(α) (or simply p_{X∗}(α)) the proper pushforward to a point of a class α ∈ A_{∗}^{T} (X). We may also write ∫_{X}(α) or deg (α) for this pushforward. Note that A_{T}^{∗} (pt) is isomorphic to A_{∗}^{T} (pt) with the opposite grading.

Let X be a T -scheme. For any mixed space X_{T} we construct a map r : A_{T}^{i} (X) → A^{i}(X_{T}). Let c ∈ A_{T}^{∗} (X). For a map Y → X_{T} and α ∈ A_{∗}(Y), we define r(c) ∩ α as follows. Let Y_{U} → Y be the pullback of the principal T - bundle X × U → X_{T}. Since Y_{U} → Y is a principal bundle, we identify A_{∗}(Y) with A_{∗}^{T} (Y_{U}). Let α_{U} ∈ A_{∗}^{T} (Y_{U}) correspond to α ∈ A_{∗}(Y). Now simply define r(c) ∩ α to be the class corresponding to c ∩ α_{U}. See [EG1, pages 620-621] for more information on the functorial properties of the map r. On the other hand, we also have a map ρ : A^{i}(X_{T}) → A_{T}^{i} (X). Indeed, let c ∈ A^{i}(X_{T}), Y → X a T -equivariant map, and β ∈ A_{∗}^{T} (Y). For any representation, there are maps Y_{T} → X_{T}. The class β is represented by a class β_{U} ∈ A_{∗+dim U−dim T} (Y_{T}) for some mixed space Y × U/T. Define ρ(c) ∩ β = c ∩ β_{U}. This is an element of A_{∗+dim U−dim T −i}(Y_{T}) ≃ A_{∗−i}^{T}(Y). Note that if X has a T -equivariant smooth envelope (e.g. X is a group embedding or char(𝕜) = 0), and V \ U has codimension more than i, then ρ and r are inverse functions; so in this case we get A_{T}^{i} (X) ≃ A^{i}(X_{T}) [EG1, Theorem 2].

Finally, for any T -scheme X, restriction to a fiber of p_{X,T} : X_{T} → U/T induces a canonical map A^{∗}(X_{T}) → A^{∗}(X). Precomposing this map with r : A_{T}^{∗} (X) → A^{∗}(X_{T}) gives a natural map ι^{∗} : A_{T}^{∗} (X) → A^{∗}(X). In general, unlike its counterpart in equivariant Chow groups, the map ι^{∗} is not surjective and its kernel is not necessarily generated in degree one, not even for toric varieties [KP]. This becomes an issue when translating results from equivariant to non-equivariant Chow cohomology. In Corollary 3.8 we give some conditions under which ι^{∗} is surjective and yields an isomorphism A_{T}^{∗} (X)_{ℚ}/ΔA_{T}^{∗} (X)_{ℚ} ≃ A^{∗}(X)_{ℚ}. Such conditions are fulfilled, among others, by complete ℚ-filtrable spherical varieties [G4].

2.2 The Białynicki-Birula decomposition

Let X be a T -scheme. Let X^{T} = ⨆_{i=1}^{m} F_{i} be the decomposition of X^{T} into connected components. A one-parameter subgroup λ : 𝔾_{m} → T is called generic if X^{𝔾_{m}} = X^{T}, where 𝔾_{m} acts on X via λ. Generic one-parameter subgroups always exist (when X is locally linearizable this certainly holds; the general case follows from this by considering the normalization of X). Now fix a generic one-parameter subgroup λ of T. For each F_{i}, we define the subset

X_{+}(F_{i}, λ) := {x ∈ X | lim_{t→0} λ(t) · x exists and is in F_{i}}.

We denote by π_{i} : X_{+}(F_{i}, λ) → F_{i} the map x ↦ lim_{t→0} λ(t)·x. Then X_{+}(F_{i}, λ) is a locally closed T -invariant subscheme of X, and π_{i} is a T -equivariant morphism. The (disjoint) union of the X_{+}(F_{i}, λ) might not cover all of X, but when it does (e.g. when X is complete), the decomposition {X_{+}(F_{i}, λ)}_{i=1}^{m} is called the Białynicki-Birula decomposition, or BB-decomposition, of X associated to λ. Each X_{+}(F_{i}, λ) is referred to as a stratum of the decomposition. If, moreover, all fixed points of the given T -action on X are isolated (i.e. X^{T} is finite), the corresponding X_{+}(F_{i}, λ) are simply called cells of the decomposition.

Definition 2.1. Let X be a T -scheme endowed with a BB-decomposition {X_{+}(F_{i}, λ)}, for some generic one-parameter subgroup λ of T. The decomposition is said to be filtrable if there exists a finite increasing sequence Σ_{0} ⊂ Σ_{1} ⊂ … ⊂ Σ_{m} of T -invariant closed subschemes of X such that:

a) Σ_{0} = ∅, Σ_{m} = X,

b) Σ_{j} \ Σ_{j−1} is a stratum of the decomposition {X_{+}(F_{i}, λ)}, for each j = 1, …, m.

In this context, it is common to say that X is filtrable. If, moreover, X^{T} is finite and the cells X_{+}(F_{i}, λ) are isomorphic to affine spaces 𝔸^{n_{i}}, then X is called T -cellular. The following result is due to Białynicki-Birula ([B1], [B2]).

Theorem 2.2. Let X be a complete T -scheme, and let λ be a generic oneparameter subgroup. If X admits an ample T -linearized invertible sheaf, then the associated BB-decomposition {X_{+}(F_{i}, λ)} is filtrable. Furthermore, if X is smooth, then X^{T} is also smooth, and for any component F_{i} of X^{T}, the map π_{i} : X_{+}(F_{i}, λ) → F_{i} makes X_{+}(F_{i}, λ) into a T -equivariant locally trivial bundle in affine spaces over F_{i}. □

Hence, smooth projective T -schemes with isolated fixed points are T -cellular.

2.3 T -linear schemes

We introduce here the main objects of our study and outline some of their relevant features.

Definition 2.3. Let T be an algebraic torus and let X be a T -scheme.

1. We say that X is T -equivariantly 0-linear if it is either empty or isomorphic to Spec (Sym(V^{∗})), where V is a finite-dimensional rational representation of T.

2. For a positive integer n, we say that X is T -equivariantly n-linear if there exists a family of T -schemes {U, Y, Z}, such that Z ⊆ Y is a T -invariant closed immersion with U its complement, Z and one of the schemes U or Y are T -equivariantly (n − 1)-linear and X is the other member of the family {U, Y, Z}.

3. We say that X is T -equivariantly linear (or simply, T -linear) if it is T - equivariantly n-linear for some n ≥ 0. T -linear varieties are varieties that are T -linear schemes.

It follows from the inductive definition that if X is T -equivariantly n-linear, then X^{H} is T -equivariantly n-linear, for any subtorus H ⊂ T. Moreover, if T → T′ is a morphism of algebraic tori, then every T′-linear scheme is also T -linear. Observe that T -linear schemes are linear schemes in the sense of Jannsen [Ja] and Totaro [To]. While Totaro’s class of linear schemes is slightly narrower than that of Jannsen, the difference is nevertheless immaterial for our purposes. In fact, one easily checks that the main result of Totaro used here, namely [To, Proposition 1], holds for the larger class. The following result is recorded in [JK].

Proposition 2.4. Let T be an algebraic torus and let T′ be a quotient of T. Let T act on T′ via the quotient map. Then the following hold:

(i) T′ is T -linear.

(ii) A T -cellular scheme is T -linear.

(iii) Every T -scheme with finitely many T -orbits is T -linear. In particular, a toric variety with dense torus T is T -linear. □

It is well-known that flag varieties, partial flag varieties and Schubert varieties come with a paving by affine spaces (due to the Bruhat decomposition), so they are all T -cellular and hence T -linear. Let B be a connected solvable linear algebraic group with maximal torus T. A result of Rosenlicht [R, Theorem 5] shows that a homogeneous space for B is isomorphic as a variety to 𝔾_{a}^{r} × 𝔾_{m}^{s}, for some r, s. As observed by Totaro [To], this implies that a B-scheme with finitely many B-orbits (e.g. a spherical variety) is linear. Nevertheless, this does not readily imply that such a scheme is T -linear, as Rosenlicht does not show that the isomorphism above may be chosen T -equivariant. Presumably his arguments can be adjusted to achieve this. In any case, we shall give a direct proof of this fact, to keep the exposition self contained.

Theorem 2.5. Let B be a connected solvable linear algebraic group with maximal torus T. Let X be a B-scheme. If B acts on X with finitely many orbits, then X is T -linear. In particular, spherical varieties are T -linear.

Proof. The following argument was shown to the author by M. Brion (personal communication). Since X is a disjoint union of B-orbits and these are T -stable, it suffices to show that every B-orbit is T -linear. Write this orbit as B/H where H is a closed subgroup of B. Let U be the unipotent radical of B. Then, we have a natural map f : B/H → B/UH and the right-hand side is a torus (for it is a homogeneous space under the torus T = B/U). Moreover, f is a Bequivariant fibration with fiber UH/H = U/(U ∩ H), which is an affine space (as it is homogeneous under U). We will show that the fibration f is T -equivariantly trivial by factoring it into T -equivariantly trivial fibrations with fiber the affine line. For this, we argue by induction on the dimension of UH/H. If this dimension is zero, then there is nothing to prove. If it is positive, then U acts non-trivially on B/H; replacing U and B with suitable quotients, we may assume that U acts faithfully. Because U is a unipotent group normalized by T, we may find a one-dimensional unipotent subgroup V of the center of U, which is normalized by T [Sp, Lemma 6.3.4]. So V is isomorphic to 𝔾_{a}, and T acts linearly on V with some weight α. By construction, V acts freely on B/H via left multiplication, and the quotient map is the natural map B/H → B/V H, which is a principal V -bundle. Since the variety B/V H is affine (e.g by the induction assumption) and V ≃ 𝔾_{a}, this bundle is trivial. The isomorphism B/H = B/V H × V, equivariant for the action of V, yields a regular function g on B/H such that g(vx) = v + g(x) for all x in B/H and v in V (identified to 𝔾_{a}). Let T act on the ring of regular functions on B/H via its action on B/H by left multiplication. We claim that g may be chosen an eigenvector of T. Indeed, write g as a sum of weight vectors g_{λ}. Then for any t in T, we obtain (tg)(vx) = g(tvx) = α(t)v + (tg)(x) which yields ∑_{λ} λ(t)g_{λ}(vx) = α(t)v + ∑_{λ} λ(t)g_{λ}(x). By viewing both sides as functions of t and using linear independence of characters, one gets g_{α}(vx) = v + g_{α}(x), and g_{λ}(vx) = g_{λ}(x) for all λ ≠ α. So these g_{λ} are invariant under V, i.e. they are regular functions on B/V H, and one may subtract them from g to get g = g_{α}. Now that the claim has been verified, it follows that the product map B/H → B/V H × V, where the second map is g_{α}, yields the desired T -equivariant isomorphism, and we conclude by induction.

2.4 Description of equivariant Chow groups

Next we state Brion’s presentation of the equivariant Chow groups of schemes with a torus action in terms of invariant cycles [Br2, Theorem 2.1]. It also shows how to recover usual Chow groups from equivariant ones.

Theorem 2.6. Let X be a T -scheme. Then the S-module A_{∗}^{T} (X) is defined by generators [Y], where Y is an invariant subvariety of X, and relations [div_{Y} (f)] − χ[Y] where f is a rational function on Y which is an eigenvector of T of weight χ. Moreover, the map A_{∗}^{T} (X) → A_{∗}(X) vanishes on ΔA_{∗}^{T} (X), and it induces an isomorphism A_{∗}^{T} (X)/ΔA_{∗}^{T} (X) → A_{∗}(X). □

Now let Γ be a connected solvable linear algebraic group with maximal torus T. If X is a Γ-scheme, then the generators of A_{∗}^{T} (X) in Theorem 2.6 can be taken to be Γ-invariant [Br2, Proposition 6.1]. In particular, if X has finitely many Γ-orbits (e.g. X is spherical), then the S-module A_{∗}^{T} (X) is finitely generated by the classes of the Γ-orbit closures. More generally, one has the following lemma.

Lemma 2.7. Let X be a T -linear scheme. Then the S-module A_{∗}^{T} (X) is finitely generated. In particular, A_{∗}(X) is a finitely generated abelian group. Moreover, if X is complete, then rational equivalence and algebraic equivalence coincide on X.

Proof. The first two assertions are easy consequences of the inductive definition of T -linear schemes (see e.g. [To]). Regarding the last one, observe that if X is complete, then the kernel of the natural morphism A_{∗}(X) → B_{∗}(X) is divisible [Fu, Example 19.1.2], and thus trivial, for A_{∗}(X) is finitely generated.

Recall that if X is a smooth equidimensional T -scheme, then A_{T}^{∗} (X) is isomorphic to the equivariant Chow group of X graded by codimension [EG1, Proposition 4].

Theorem 2.8 ([Br2], [VV]). Let X be a smooth T -variety. If X is complete, then the S_{ℚ}-module A_{∗}^{T} (X)_{ℚ} is free. Moreover, the restriction homomorphism i_{T}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (X^{T})_{ℚ} is injective, and its image is the intersection of all the images of the restriction homomorphisms i_{T,H}^{∗} : A_{T}^{∗} (X^{H})_{ℚ} → A_{T}^{∗} (X^{T})_{ℚ}, where H runs over all subtori of codimension one in T.

A few comments are in order here. First, Brion showed that Theorem 2.8 holds in the special case that X is projective [Br2, Theorems 3.2 and 3.3]. Later on, Vezzosi and Vistoli [VV] generalized Brion’s results to the setting of equivariant higher K-theory and established the corresponding analogue of Theorem 2.8, which holds for X complete [VV, Corollary 5.11]. From this, by making appropriate changes in the proofs of [VV, Proposition 5.13 and Theorem 5.4], one obtains Theorem 2.8 in its full form. The details can be found in a preprint version of [VV] (arXiv version 3). Alternatively, see [Kr, Sections 9 and 10], where the results of [VV, Section 5] have been generalized to equivariant higher Chow groups.

In characteristic zero Theorem 2.8 extends to all possibly singular complete varieties [G3, Section 7]. See the Appendix for a review of the main results in this regard.

The next lemma will become relevant later, when integrality of the equivariant operational Chow rings is discussed (cf. Lemma 3.3). It is essentially due to Brion, del Baño and Karpenko.

Lemma 2.9. Let X be a smooth projective T -variety. Then the following are equivalent.

(i) A_{∗}(X^{T}) is ℤ-free.

(ii) A_{∗}^{T} (X) is S-free.

(iii) A_{∗}(X) is ℤ-free.

If moreover X is T -linear, then any (and hence all) of these conditions hold.

Proof. The implication (i)⇒(ii) follows from [Br2, Corollary 3.2.1], as any smooth projective variety is filtrable (Theorem 2.2). That (ii) implies (iii) is a consequence of Theorem 2.6. To show that (iii) implies (i) we use a result of del Baño [dB, Theorem 2.4] and Karpenko [Ka, Section 6]. Namely, let λ be a generic one-parameter subgroup of T, and let X^{T} = ⨆_{i} F_{i} be the decomposition of X^{T} into connected components. Then, for every non-negative integer j ≤ dim (X), there is a natural isomorphism \xymatrix{⨁_{i} A^{j−d_{i}}(F_{i}) \ar[r]^{≃} & A^{j}(X)} , where d_{i} is the codimension of X_{+}(F_{i}, λ) in X (all spaces involved are smooth, so there is an intersection product on the Chow groups graded by codimension). These isomorphisms yield the assertion (iii)⇒(i). Finally, if X is a smooth projective T -linear variety, then, in particular, it is a projective linear variety, and so it satisfies the Künneth formula (see below). Now Theorem 2.10 (ii) implies that condition (iii) of the lemma holds for X. This concludes the argument.

For any schemes X and Y, one has a Künneth map

A_{∗}(X) ⊗ A_{∗}(Y) → A_{∗}(X × Y),

taking [V] ⊗ [W] to [V × W], where V and W are subvarieties of X and Y. This is an isomorphism only for very special schemes, e.g. linear schemes [To, Proposition 1]; but when it is, strong consequences can be derived from it, as we shall see below. Let us start with the following result due to Ellingsrud and Stromme [ES, Theorem 2.1].

Theorem 2.10. Let X be a smooth complete variety. Assume that the rational equivalence class δ of the diagonal Δ(X) ⊆ X × X is in the image of the Künneth map A_{∗}(X) ⊗ A_{∗}(X) → A_{∗}(X × X). Let δ = ∑ u_{i} ⊗ v_{i} be a corresponding decomposition of δ, where u_{i}, v_{i} ∈ A_{∗}(X). Then

(i) The v_{i} generate A_{∗}(X), i.e. any z ∈ A_{∗}(X) has the form ∑(u_{i} · z)v_{i}.

(ii) Numerical and rational equivalence coincide on X. In particular, A_{∗}(X) is a free ℤ-module.

(iii) If k = ℂ, then the cycle map cl_{X} : A_{∗}(X) → H_{∗}(X, ℤ) is an isomorphism. In particular, the homology and cohomology groups of X vanish in odd degrees.

Now consider a smooth complex algebraic variety X with an action of a complex algebraic torus T. Together with a cycle map cl_{X} : A^{∗}(X) → H^{∗}(X, ℤ) (which doubles degrees [Fu, Corollary 19.2]), there is also an equivariant cycle map cl_{X}^{T} : A_{T}^{∗} (X) → H_{T}^{∗} (X, ℤ) where H_{T}^{∗} (X, ℤ) denotes equivariant cohomology with integral coefficients, see [EG1, Section 2.8]. Next is a version of Theorem 2.10 (iii) for cl_{X}^{T}.

Proposition 2.11. Let X be a smooth complete complex T -variety. If the class of the diagonal Δ(X) ⊆ X × X is in the image of the Künneth map A_{∗}(X) ⊗ A_{∗}(X) → A_{∗}(X × X), then the equivariant cycle map

cl_{X}^{T} : A_{T}^{∗} (X) → H_{T}^{∗} (X, ℤ)

is an isomorphism. In particular, this holds if X is T -linear.

Proof. In view of Theorem 2.10 (iii), the given hypothesis on X imply that cl_{X} is an isomorphism, and that X has no integral cohomology in odd degrees. Then the spectral sequence associated to the fibration X ×_{T} ET → BT collapses, where ET → BT is the universal T -bundle. So the S-module H_{T}^{∗} (X, ℤ) is free, and the map H_{T}^{∗} (X, ℤ)/ΔH_{T}^{∗} (X, ℤ) → H^{∗}(X, ℤ) is an isomorphism. These results, together with the graded Nakayama Lemma, yield surjectivity of the equivariant cycle map cl_{X}^{T} : A_{T}^{∗} (X) → H_{T}^{∗} (X, ℤ). To show injectivity, we proceed as follows. First, choose a basis z_{1}, .., z_{n} of H^{∗}(X, ℤ). Now identify that basis with a basis of A^{∗}(X) (via cl_{X}) and lift it to a generating system of the S-module A_{T}^{∗} (X). Then this generating system is a basis, since its image under the equivariant cycle map is a basis of H_{T}^{∗} (X, ℤ). For the last assertion of the proposition, simply recall that if X is T -linear, then the Künneth map is an isomorphism [To, Proposition 1].

2.5 Equivariant Localization

Let T be an algebraic torus. The following is the localization theorem for equivariant Chow groups [Br2, Corollary 2.3.2].

Theorem 2.12. Let X be a T -scheme. If X is locally linearizable, then the S-linear map i_{T ∗} : A_{∗}^{T} (X^{T}) → A_{∗}^{T} (X) is an isomorphism after inverting all non-zero elements of Δ. □

For later use, we prove a slightly more general statement.

Proposition 2.13. Let X be a T -scheme, and let H ⊂ T be a closed subgroup. Then the S-linear map i_{H∗} : A_{∗}^{T} (X^{H}) → A_{∗}^{T} (X) becomes an isomorphism after inverting finitely many characters of T that restrict non-trivially to H.

Before proving this proposition, we need a technical lemma. We would like to thank M. Brion for suggesting a simplified proof of this fact.

Lemma 2.14. Let X be an affine T -scheme. Let H be a closed subgroup of T. Then the ideal of the fixed point subscheme X^{H} is generated by all regular functions on X which are eigenvectors of T with a weight that restricts nontrivially to H.

Proof. Recall that X^{H} is the largest closed subscheme of X on which H acts trivially. In other words, the ideal I of X^{H} is the smallest H-stable ideal of k[X] such that H acts trivially on the quotient k[X]/I. So I is T -stable and hence the direct sum of its T -eigenspaces. Moreover, if f ∈ k[X] is a T -eigenvector of weight χ which restricts non-trivially to H, then f ∈ I. Indeed, let \overline{f} be the image of f in k[X]/I. Notice that \overline{f} is a T -eigenvector of the same weight χ as f. Since H acts trivially on k[X]/I, we obtain the identity \overline{f} = h · \overline{f} = χ(h)\overline{f}, valid for all h ∈ H. Nevertheless, there exists h_{0} ∈ H such that χ(h_{0}) ≠ 1, by our assumption on χ. Substituting this information into the above identity yields \overline{f} = 0, equivalently, f ∈ I. Thus, I contains the ideal J generated by all such functions f. But k[X]/J is a trivial H-module by construction, and hence I = J by minimality.

Proof of Proposition 2.13. In virtue of Lemma 2.14, the proof is an easy adaptation of Brion’s proof of [Br2, Corollary 2.3.2], so we provide only a sketch of the crucial points. First, assume that X is locally linearizable, i.e. X is a finite union of T -stable affine open subsets X_{i}. Lemma 2.14 implies that the ideal of each fixed point subscheme X_{i}^{H} is generated by all regular functions on X_{i} which are eigenvectors of T with a weight that restricts non-trivially to H. Choose a finite set of such generators (f_{ij}), with respective weights χ_{ij}. From Theorem 2.6 we know that the S-module A_{∗}^{T} (X) is generated by the classes of T -invariant subvarieties of X. Now let Y ⊂ X be a T -invariant subvariety of positive dimension. If Y is not fixed pointwise by H, then one of the f_{ij} defines a non-zero rational function on Y. Then, in the Chow group, we have χ_{ij}[Y] = [div_{Y} f_{ij}]. So after inverting χ_{ij}, we get [Y] = χ_{ij}^{−1}[div_{Y} f_{ij}]. Arguing by induction on the dimension of Y, we obtain that i_{∗} becomes surjective after inverting the χ_{ij}’s. A similar argument, using these χ_{ij}’s in the proof of [Br2, Corollary 2.3.2], shows that i^{∗} is injective after localization.

Finally, if X is not locally linearizable, choose an equivariant birational envelope π : X̃ → X, where X̃ is normal (and possibly not irreducible). Let U ⊂ X be the open subset where π is an isomorphism. Set Z = X \ U and E = π^{−1}(Z). Then, by [FMSS, Lemma 2] and [EG2, Lemma 7.2], there is a commutative diagram

\xymatrix{A_{∗}^{T} (E^{H}) \ar[r] \ar[d]^{i_{H∗}} & A_{∗}^{T} (Z^{H}) ⊕ A_{∗}^{T} (X̃^{H}) \ar[r] \ar[d]^{i_{H∗}} & A_{∗}^{T} (X^{H}) \ar[r] \ar[d]^{i_{H∗}} & 0 \\ A_{∗}^{T} (E) \ar[r] & A_{∗}^{T} (Z) ⊕ A_{∗}^{T} (X̃) \ar[r] & A_{∗}^{T} (X) \ar[r] & 0.}

Observe that E and Z have strictly smaller dimension than X. Moreover, E and X̃ are locally linearizable. Applying Noetherian induction and the previous part of the proof, we get that the first two left vertical maps become isomorphisms after localization; hence so does the third one.

3 Equivariant Kronecker duality and Localization

3.1 Equivariant Kronecker duality schemes

Definition 3.1. Let X be a complete T -scheme. We say that X satisfies T -equivariant Kronecker duality if the following conditions hold:

(i) A_{∗}^{T} (X) is a finitely generated S-module.

(ii) The equivariant Kronecker duality map

K_{T} : A_{T}^{∗} (X) ⟶ Hom_{S}(A_{∗}^{T} (X), S) α ↦ (β ↦ p_{X∗}(β ∩ α))

is an isomorphism of S-modules.

Likewise, we say that X satisfies rational T -equivariant Kronecker duality if the S_{ℚ}-modules A_{∗}^{T} (X)_{ℚ} and A_{T}^{∗} (X)_{ℚ} satisfy the conditions (i) and (ii) with rational coefficients. Remark 3.2. The equivariant Kronecker duality map is functorial for morphisms between complete T -schemes. Indeed, let f : X̃ → X be an equivariant (proper) morphism of complete T -schemes. For any ξ ∈ A_{T}^{∗} (X), we have

∫X̃ f^{∗}(ξ) ∩ z = ∫_{X} f_{∗}(f^{∗}(ξ) ∩ z) = ∫_{X}(ξ ∩ f_{∗}(z)),

due to the projection formula [Fu]. This identity implies the commutativity of the diagram

\xymatrix{A_{T}^{∗} (X) \ar[r]^{f^{∗}} \ar[d]_{K_{T}} & A_{T}^{∗} (X̃) \ar[d]^{K_{T}} \\ Hom_{S}(A_{∗}^{T} (X), S) \ar[r]^{(f_{∗})^{t}} & Hom_{S}(A_{∗}^{T} (X̃), S),}

where (f_{∗})^{t} is the transpose of f_{∗} : A_{∗}^{T} (X̃) → A_{∗}^{T} (X).

It follows from Definition 3.1 that if X satisfies T -equivariant Kronecker duality, then the S-module A_{T}^{∗} (X) is finitely generated and torsion free. In particular, if T is one dimensional, i.e. T = 𝔾_{m}, then A_{𝔾_{m}}^{∗}(X) is a finitely generated free module over A_{𝔾_{m}}^{∗} = ℤ[t]. Moreover, if X is projective and smooth, then A_{∗}(X^{𝔾_{m}}) is a finitely generated free abelian group (Lemma 2.9). As it stems from the previous paragraph, not all smooth varieties with a torus action satisfy Equivariant Kronecker duality. For a more concrete example, consider the trivial action of T on a projective smooth curve. In this case, one checks that K_{T} is an extension of the non-equivariant Kronecker duality map K. But, as pointed out in [FMSS], the kernel of K in degree one is the Jacobian of the curve, which is non-trivial if the curve has positive genus.

Lemma 3.3. Let X be a smooth complete T -variety. Then X satisfies rational T -equivariant Kronecker duality if and only if it satisfies the rational non-equivariant Kronecker duality, i.e. K : A^{i}(X)_{ℚ} → Hom(A_{i}(X), ℚ) is an isomorphism for all i. If, moreover, X is projective and A_{∗}(X^{T}) is ℤ-free, then the equivalence holds over the integers.

Proof. Both assertions are proved similarly, so we focus on the second one. Since X is smooth and projective, the assumption on A_{∗}(X^{T}) implies that A_{∗}^{T} (X) is a free S-module (Lemma 2.9; cf. Theorem 2.8). Now, by Poincaré duality [EG1, Proposition 4], A_{T}^{∗} (X) is isomorphic to A_{∗}^{T} (X); so it is also a free S-module. By the graded Nakayama lemma, K_{T} is an isomorphism if and only if

\overline{K_{T}} : A_{T}^{∗} (X)/ΔA_{T}^{∗} (X) → Hom_{S}(A_{∗}^{T} (X), S)/ΔHom_{S}(A_{∗}^{T} (X), S)

is an isomorphism. But freeness of A_{∗}^{T} (X) yields an isomorphism

Hom_{S}(A_{∗}^{T} (X), S)/ΔHom_{S}(A_{∗}^{T} (X), S) ≃ Hom(A_{∗}^{T} (X)/ΔA_{∗}^{T} (X), ℤ)

and the later identifies to Hom(A_{∗}(X), ℤ), by Theorem 2.6. On the other hand, by Theorem 2.6 again, the map A_{T}^{∗} (X)/ΔA_{T}^{∗} (X) → A^{∗}(X) is an isomorphism. These facts, together with the commutativity of the diagram below

\xymatrix{A_{T}^{∗} (X) \ar[r]^{K_{T}} \ar[d] & Hom_{S}(A_{∗}^{T} (X), S) \ar[d] \\ A^{∗}(X) \ar[r]^{K} & Hom(A_{∗}(X), ℤ),}

yield the content of the lemma.

Next we show that complete T -linear schemes satisfy equivariant Kronecker duality. For this, the main ingredient is the following result, due to Totaro [To] in the non-equivariant case. Recall that a T -scheme X is said to satisfy the equivariant Künneth formula if the Künneth map (or exterior product [EG1])

A_{∗}^{T} (X) ⊗_{S} A_{∗}^{T} (Y) → A_{∗}^{T} (X × Y)

is an isomorphism for any T -scheme Y.

Proposition 3.4. Let X be a T -scheme. If X is T -linear, then it satisfies the equivariant Künneth formula.

Proof. When X is T -linear, one can choose representations V of T so that X_{T} is linear, see e.g. [Br2, Section 2.2] and [P, Section 1]. Now the result follows from [To, Proposition 1].

Proposition 3.5. If X is a complete T -linear scheme, then the equivariant Kronecker map

K_{T} : A_{T}^{∗} (X) → Hom_{S}(A_{∗}^{T} (X), S)

is an isomorphism.

This result follows quite formally from Proposition 3.4, as in the nonequivariant case [FMSS, Theorem 3], so we only sketch the proof. To define the inverse to K_{T}, given a S-module homomorphism φ : A_{∗}^{T} (X) → S, we construct an element c_{φ} ∈ A_{T}^{∗} (X). Since the S-module A_{∗}^{T} (X) is finitely generated, we can assume, without loss of generality, that φ is homogeneous [Bo, Part II, Section 11.6]. Bearing this in mind, given a homomorphism φ : A_{∗}^{T} (X) → S of degree −λ, we build c_{φ} ∈ A_{T}^{λ} (X) as follows. For a T -map f : Y → X, the corresponding homomorphism f^{∗}c_{φ} := c_{φ}(f) : A_{∗}^{T} (Y) → A_{∗−λ}^{T}(Y) is defined to be the composite

\xymatrix{A_{∗}^{T} (Y) \ar[r]^{(γ_{f})_{∗}} & A_{∗}^{T} (X × Y) \ar[r]^{≃} & A_{∗}^{T} (X) ⊗_{S} A_{∗}^{T} (Y) \ar[r]^{φ⊗id} & S ⊗_{S} A_{∗}^{T} (Y) ≃ A_{∗}^{T} (Y)_{,}}

where (γ_{f})_{∗} denotes the proper pushforward along the graph of f, and the second displayed map is the Künneth isomorphism (Proposition 3.4). The verification that c_{φ} satisfies the compatibility axioms, and that this construction indeed gives the inverse to K_{T}, is the same as in [FMSS].

For c ∈ A_{T}^{λ} (X) and z ∈ A_{∗}^{T} (X), we write c(z) for deg(c ∩ z). The next two corollaries describe the cap and cup product structures. They are easily deduced from the previous proposition, cf. [FMSS, Corollaries 1 and 2].

Corollary 3.6. Let f : Y → X, c ∈ A_{T}^{λ} (X), z ∈ A_{m}^{T}(Y). Suppose (γ_{f})_{∗}(z) = ∑ u_{i} ⊗ v_{i} with u_{i} ∈ A_{p(i)}^{T}(X) and v_{i} ∈ A_{m−p(i)}^{T}(Y). Then f^{∗}c ∩ z = ∑_{p(i)≤λ} c(u_{i})v_{i}.

Corollary 3.7. Let c ∈ A_{T}^{λ} (X), c′ ∈ A_{T}^{μ} (X), and z ∈ A_{m}^{T}(X), where m ≤ λ + μ. Write δ_{∗}(z) = ∑ u_{i} ⊗ v_{i} with u_{i} ∈ A_{p(i)}^{T}(X) and v_{i} ∈ A_{m−p(i)}^{T}(X). Then (c ∪ c′)(z) = ∑_{m−μ≤p(i)≤λ} c(u_{i})c′(v_{i}).

For a T -scheme X, there is a natural map ι^{∗} : A_{T}^{∗} (X) → A^{∗}(X) (Section 2.1). In general, when X is singular, ι^{∗} may not be surjective, and its kernel may not be generated in degree one [KP]. Next we describe a class of possibly singular T -schemes for which the map ι^{∗} is well-behaved. This yields the compatibility of our product formulas with those of [FMSS].

Corollary 3.8. Let X be a complete T -scheme. If X is T -linear and A_{∗}^{T} (X) is S-free, then the map A_{T}^{∗} (X)/ΔA_{T}^{∗} (X) → A^{∗}(X), induced by ι^{∗}, is an isomorphism.

Proof. Proposition 3.5 together with freeness of A_{∗}^{T} (X) yield

A_{T}^{∗} (X)/ΔA_{T}^{∗} (X) ≃ Hom_{S}(A_{∗}^{T} (X), S)/ΔHom_{S}(A_{∗}^{T} (X), S) ≃  Hom_{ℤ}(A_{∗}^{T} (X)/ΔA_{∗}^{T} (X), ℤ).

Furthermore, by Theorem 2.6, the term on the right hand side corresponds to Hom(A_{∗}(X), ℤ), which, in turn, is isomorphic to A^{∗}(X), due to the nonequivariant version of Kronecker duality [To, Proposition 1]. Considering this information alongside the commutative diagram

\xymatrix{A_{T}^{∗} (X) \ar[r]^{K_{T}} \ar[d] & Hom_{S}(A_{∗}^{T} (X), S) \ar[d] \\ A^{∗}(X) \ar[r]^{K} & Hom(A_{∗}(X), ℤ),}

produces the content of the corollary.

The conditions of Corollary 3.8 are satisfied by possibly singular T -cellular varieties (e.g. Schubert varieties). With ℚ-coefficients, the corresponding statement is satisfied by ℚ-filtrable spherical varieties [G4]. This class includes all rationally smooth projective equivariant embeddings of reductive groups [G4].

3.2 Localization for T -equivariant Kronecker duality schemes

From the viewpoint of algebraic torus actions, the main attribute of equivariant Kronecker duality schemes is that they supply a somewhat more intrinsic background for establishing localization theorems on integral equivariant Chow cohomology.

Theorem 3.9. Let X be a complete T -scheme satisfying T -equivariant Kronecker duality. Let H ⊂ T be a subtorus of T and let i_{H} : X^{H} → X be the inclusion of the fixed point subscheme. If X^{H} also satisfies T -equivariant Kronecker duality, then the morphism

i_{H}^{∗} : A_{T}^{∗} (X) → A_{T}^{∗} (X^{H})

becomes an isomorphism after inverting finitely many characters of T that restrict non-trivially to H. In particular, i_{H}^{∗} is injective over ℤ.

Proof. By Proposition 2.13, the localized map (i_{H}∗)_{F} : A_{∗}^{T} (X^{H})_{F} → A_{∗}^{T} (X)_{F} is an isomorphism, where F is a finite family of characters of T that restrict non-trivially to H. Now consider the commutative diagram

\xymatrix{A_{T}^{∗} (X) \ar[r]^{i_{H}^{∗}} \ar[d] & A_{T}^{∗} (X^{H}) \ar[d] \\ _{Hom_{S} (A_{∗}^{T} (X), S)} \ar[r]^{(i_{H∗})^{t}} & Hom_{S} (A_{∗}^{T} (X^{H}), S),}

where (i_{H∗})^{t} represents the transpose of i_{H∗} : A_{∗}^{T} (X^{H}) → A_{∗}^{T} (X) (commutativity follows from Remark 3.2, because i_{H} is proper). By our assumptions on X and X^{H}, both vertical maps are isomorphisms. Moreover, after localization at F, the above commutative diagram becomes

\xymatrix{A_{T}^{∗} (X)_{F} \ar[r]^{(i_{H}^{∗})_{F}} \ar[d] & A_{T}^{∗} (X^{H})_{F} \ar[d] \\ (Hom_{S} (A_{∗}^{T} (X), S))_{F} \ar[r]^{((i_{H∗})^{t})_{F}} & (Hom_{S}(A_{∗}^{T} (X^{H}), S))_{F}.}

Since A_{∗}^{T} (X) is a finitely generated S-module (as X satisfies equivariant Kronecker duality), localization commutes with formation of Hom (see [Ei, Prop. 2.10, p. 69]), and so

A_{T}^{∗} (X)_{F} ≃ (Hom_{S} (A_{∗}^{T} (X), S))_{F} ≃ Hom_{S_{F}} (A_{∗}^{T} (X)_{F}, S_{F}).

Similarly, for X^{H} we obtain

A_{T}^{∗} (X^{H})_{F} ≃ (Hom_{S}(A_{∗}^{T} (X^{H}), S))_{F} ≃ Hom_{S_{F}} (A_{∗}^{T} (X^{H})_{F}, S_{F}).

In other words, the bottom map from the preceding diagram fits in the commutative square

\xymatrix{(Hom_{S} (A_{∗}^{T} (X), S))_{F} \ar[r]^{((i_{H∗})^{t})_{F}} \ar[d] & (Hom_{S}(A_{∗}^{T} (X^{H}), S))_{F} \ar[d] \\ Hom_{S_{F}} (A_{∗}^{T} (X)_{F}, S_{F}) \ar[r]^{((i_{H∗})_{F})^{t}} & Hom_{S_{F}} (A_{∗}^{T} (X^{H})_{F}, S_{F}),}

where the vertical maps are natural isomorphisms. But we already know that (i_{H∗})_{F} is an isomorphism, hence so are ((i_{H∗})_{F})^{t}, ((i_{H∗})^{t})_{F} and (i_{H}^{∗})_{F}.

Finally, to prove the last assertion of the theorem, recall that the S-module A_{T}^{∗} (X) is finitely generated and torsion free (Definition 3.1). Hence the natural map A_{T}^{∗} (X) → A_{T}^{∗} (X) ⊗_{S} Q is injective, where Q is the quotient field of S. In particular, the (also natural) map A_{T}^{∗} (X) → A_{T}^{∗} (X)_{F} is injective. This, together with the first part of the theorem, yields injectivity of i_{H}^{∗}. We are done.

Corollary 3.10. Let X be a complete T -scheme. Let H be a codimension-one subtorus of T. If X is T -linear, then the pullback i_{H}^{∗} : A_{T}^{∗} (X) → A_{T}^{∗} (X^{H}) is injective over ℤ.

Proof. If X is T -linear, then so is X^{H}. Now use Proposition 3.5 and Theorem 3.9.

Let X be a complete T -linear scheme. It follows from Corollary 3.10 that the image of the injective map i_{T}^{∗} : A_{T}^{∗} (X) → A_{T}^{∗} (X^{T}) is contained in the intersection of the images of the (also injective) maps i_{T,H}^{∗} : A_{T}^{∗} (X^{H}) → A_{T}^{∗} (X^{T}), where H runs over all subtori of codimension one in T. When the image of i_{T}^{∗} is exactly the intersection of the images of the maps i_{T,H}^{∗} we say, following [G3], that X has the Chang-Skjelbred property (or CS property). If the defining condition holds over ℚ rather than ℤ, we say that X has the rational CS property. By Theorem 2.8, any smooth complete T -scheme has the rational CS property; by Theorem A.6, so does any complete T -scheme in characteristic zero. It would be interesting to determine, in arbitrary characteristic, which complete, possibly singular, T -linear schemes satisfy the CS property. For instance, toric varieties are known to have this property [P]. We anticipate that this also holds for projective embeddings of semisimple groups of adjoint type (this shall be pursued elsewhere). For the corresponding problem with rational coefficients, we provide an answer next.

Theorem 3.11. Let X be a complete T -linear scheme. If there exists an equivariant envelope π : X̃ → X with X̃ smooth, then X has the rational CS property. In particular, projective embeddings of connected reductive linear algebraic groups have the rational CS property in arbitrary characteristic.

Proof. Let u ∈ A_{T}^{∗} (X^{T})_{ℚ} be such that u ∈ ⋂_{H⊂T} Im(i_{T,H}^{∗}), where the intersection runs over all codimension-one subtori H of T. Our task is to show that u ∈ Im(i_{T}^{∗})_{ℚ}. First, observe that there is a commutative diagram

\xymatrix{ & A_{T}^{∗} (X)_{ℚ} \ar[rr]^{π^{∗}} \ar[d]^{i_{T}^{∗}} \ar[ddl]_{i_{H}^{∗}} &  & A_{T}^{∗} (X̃)_{ℚ} \ar[d]^{ĩ_{T}^{∗}} \ar[ddl]_{ĩ_{H}^{∗}} \\  & A_{T}^{∗} (X^{T})_{ℚ} \ar[rr]^{_{π_{T} _{∗}}} &  & A_{T}^{∗} (X̃^{T})_{ℚ} \\ A_{T}^{∗} (X^{H})_{ℚ} \ar[rr]^{_{π_{H}_{∗}}} \ar@{-->}[ur]_{i_{T,H}^{∗}} &  & A_{T}^{∗} (X̃^{H})_{ℚ} \ar@{-->}[ur]_{ĩ_{T,H}^{∗}}}

obtained by combining and comparing the sequences that [EG2, Lemma 7.2] and Theorem A.1 assign to the envelopes π : X̃ → X, π_{H} : X̃^{H} → X^{H} and p_{T} : X̃^{T} → X^{T} (cf. [G3, proof of Theorem 4.4]. From the diagram it follows that π_{T}^{∗}(u) is in the image of ĩ_{T,H}^{∗}. Hence, π_{T}^{∗}(u) is in the intersection of the images of all ĩ_{T,H}^{∗}, where H runs over all codimension-one subtori of T. Since X̃ is known to have the rational CS property (Theorem 2.8), π_{T}^{∗}(u) is in the image of ĩ_{T}^{∗}. So let y ∈ A_{T}^{∗} (X̃) be such that ĩ_{T}^{∗} (y) = π_{T}^{∗} (u). To conclude the proof, we need to check that y is in the image of π^{∗}. In view of equivariant Kronecker duality (Proposition 3.5), this is equivalent to checking that the dual of y, namely, φ̃ := K_{T} (y) is in the image of π_{∗}^{t}, the transpose of the surjective morphism π_{∗} : A_{∗}^{T} (X̃)_{ℚ} → A_{∗}^{T} (X)_{ℚ}. Also, we should observe that the functor K_{T} (−) transforms the previous commutative diagram into another one involving the corresponding dual modules Hom(A_{∗}^{T} (−)_{ℚ}, S_{ℚ}). Now set φ_{u} := K_{T} (u). By construction, for every codimension-one subtorus H, there exists φ_{u}^{H} : A_{∗}^{T} (X^{H})_{ℚ} → S_{ℚ} such that φ_{u} = φ_{u}^{H} ∘ i_{T,H}_{∗}. In fact, we can place this information into a commutative diagram:

\xymatrix{A_{∗}^{T} (X̃^{T})_{ℚ} \ar@{->>}[d]^{π_{∗}^{T}} \ar[r]^{^{ĩ_{T,H}^{∗}}} & A_{∗}^{T} (X̃^{H})_{ℚ} \ar@{->>}[d]^{π_{∗}^{H}} \ar[r] & A_{∗}^{T} (X̃)_{ℚ} \ar@{->>}[d]^{π_{∗}} \ar[ddl]_{φ̃} \\ A_{∗}^{T} (X^{T})_{ℚ} \ar[r]^{_{i_{T,H}_{∗}}} \ar[dr]^{φ_{u}} & A_{∗}^{T} (X^{H})_{ℚ} \ar[r] \ar[d]^{φ_{u}^{H}} & A_{∗}^{T} (X)_{ℚ} \ar@{-->}[dl]^{φ (∃?)} \\  & S_{ℚ}.}

In this form, our task reduces to showing that there exists φ making the dotted arrow into a solid arrow. Bearing this in mind, we claim that φ̃ is zero on the kernel of π_{∗}. Indeed, let v ∈ A_{∗}^{T} (X̃)_{ℚ} be such that π_{∗}(v) = 0. By the localization theorem there exists a product of non-trivial characters χ_{1} ⋯ χ_{m} such that χ_{1} ⋯ χ_{m} · v is in the image of ĩ_{∗}^{T} : A_{∗}^{T} (X̃^{T})_{ℚ} → A_{∗}^{T} (X̃)_{ℚ}. As both of these S-modules are free, then, unless v is zero, we have χ_{1} ⋯ χ_{m} · v ≠ 0. Let w be such that ĩ_{∗}^{T} (w) = χ_{1} · χ_{m} · v. By commutativity of the diagram, i_{∗}^{T} (π_{∗}^{T} (w)) = 0. But i_{∗}^{T} is injective by Theorem 3.9, so π_{∗}^{T} (w) = 0. Thus

φ̃(χ_{1} ⋯ χ_{m} · v) = χ_{1} ⋯ χ_{m} · φ̃(v) = φ_{u}(π_{∗}^{T} (w)) = 0.

As S_{ℚ} has no torsion, we get φ̃(v) = 0, which proves the claim. Using this, one easily defines φ with the sought-after properties. Finally, for the last assertion of the theorem, recall that group embeddings are a special class of spherical varieties known to have resolutions of singularities in arbitrary characteristic [BK, Chapter 6].

Remark 3.12. Theorem 3.11 and its proof can be readily translated into the language of equivariant operational K-theory with ℤ-coefficients. See [G3, Section 6] for a presentation of the (integral) equivariant operational K-theory of projective group embeddings.

4 Rational equivariant Chow cohomology of spherical varieties

Throughout this section we work in characteristic zero. The aim is to describe the rational equivariant Chow cohomology of a spherical variety by comparing it with that of an equivariant resolution, using the rational CS property (Theorem 3.11) and equivariant Kronecker duality (Proposition 3.5). The main result (Theorem 4.8), inspired by [Br2, Theorem 7.3], is an extension of Brion’s description to the setting of equivariant operational Chow groups.

In what follows, we denote by G a connected reductive linear algebraic group with Borel subgroup B and maximal torus T ⊂ B. We denote by W the Weyl group of (G, T). Observe that W is generated by reflections {s_{α}}_{α∈Φ}, where Φ stands for the set of roots of (G, T). Recall that S_{ℚ}^{W} = (A_{T}^{∗}_{ℚ})^{W} = A_{G}^{∗} ⊗ ℚ. For the purposes of this section, we shall assume that G-spherical varieties are locally linearizable for the induced T -action.

4.1 Preliminaries

Recall that any spherical G-variety contains only finitely many G-orbits; as a consequence, it contains only finitely many fixed points of T. Moreover, since char(𝕜) = 0, any spherical G-variety X admits an equivariant resolution of singularities, i.e., there exists a smooth G-variety X̃ together with a proper birational G-equivariant morphism π : X̃ → X. Then the G-variety X̃ is also spherical; if moreover X is complete, we may arrange so that X̃ is projective. Notice that, in general, a resolution of singularities need not be an equivariant envelope. The next result gives an important class of spherical varieties for which equivariant resolutions are equivariant envelopes. We thank M. Brion for leading us to the following proof.

Proposition 4.1. Let X be a normal simply-connected spherical G-variety (i.e. the B-isotropy group of its dense orbit is connected). Let f : X̃ → X be a proper birational morphism. Then f is an equivariant envelope.

Proof. It suffices to show that every B-orbit in X is the isomorphic image via f of a B-orbit in X̃. So let O = (B) · x = B/B_{x} be an orbit in X. It follows from [BJ1] that O has a connected isotropy group. The preimage f^{−1}(O) ⊂ X̃ is of the form B ×^{B_{x}} F, where F denotes the fiber p^{−1}(x). Since F is connected and complete (by Zariski’s main theorem), it contains a fixed point y of the connected solvable group B_{x}. Then the orbit B·y in X̃ is mapped isomorphically to B · x.

Remark 4.2. Examples of simply-connected spherical varieties include (normal) G × G-equivariant embeddings of G.

Now we record a few notions and results from [Br2, Section 7] needed in our task. A subtorus H ⊂ T is called regular if its centralizer C_{G}(H) is equal to T; otherwise H is called singular. A subtorus of codimension one is singular if and only if it is the kernel of some positive root α. In this case, α is unique and the group C_{G}(H) is the product of H with a subgroup Γ isomorphic to SL_{2} or P SL_{2}. Furthermore, if X is a G-variety, then X^{H} inherits an action of C_{G}(H)/H, a quotient of Γ.

Proposition 4.3 ([Br2, Proposition 7.1]). Let X be a spherical G-variety. Let H ⊂ T be a subtorus of codimension one. Then each irreducible component of X^{H} is a spherical C_{G}(H)-variety. Moreover,

(i) If H is regular, then X^{H} is at most one-dimensional.

(ii) If H is singular, then X^{H} is at most two-dimensional. If moreover X is complete and smooth, then any two-dimensional connected component of X^{H} is (up to a finite, purely inseparable equivariant morphism) either a rational ruled surface 𝔽_{n} = ℙ(O_{ℙ^{1}} ⊕ O_{ℙ^{1}}(n)), where C_{G}(H) acts through the natural action of SL_{2}, or the projective plane, where C_{G}(H) acts through the projectivization of a non-trivial SL_{2}-module of dimension three. □

For later use, we record Brion’s presentation of the rational equivariant Chow rings of nonsingular ruled surfaces. We follow closely the notation and conventions of [Br2, Section 7]. Let D be the torus of diagonal matrices in SL_{2}, and let α be the character of D given by

α (\begin{matrix}t & 0 \\ 0 & t^{−1} \end{matrix}) = t^{2}.

We identify the character ring of D with ℚ[α]. Now consider a rational ruled surface 𝔽_{n} with ruling π : 𝔽_{n} → ℙ^{1}. Notice that 𝔽_{n} has exactly four fixed points x, y, z, t of D, where x, y (resp. z, t) are mapped to 0 (resp. ∞) by π.

Moreover, we may assume that x and z lie in one G-invariant section of π, and that y and t lie in the other G-invariant section. With this ordering of the fixed points, we identify A_{T}^{∗} (𝔽_{n}^{D})_{ℚ} with ℚ[α]^{4}. In contrast, denote by ℙ(V) the projectivization of a nontrivial SL_{2}-module V of dimension three. The weights of D in V are either −2α, 0, 2α (in the case when V = sl_{2}) or −α, 0, α (in the case when V = k^{2} ⊕ k). We denote by x, y, z the corresponding fixed points of D in ℙ(V), and we identify A_{T}^{∗} (ℙ(V)^{D})_{ℚ} with ℚ[α]^{3}.

Proposition 4.4. Notation being as above, the image of

i_{D}^{∗} : A_{D}^{∗}(𝔽_{n})_{ℚ} → ℚ[α]^{4}

consists of all (f_{x}, f_{y}, f_{z}, f_{t}) such that f_{x} ≡ f_{y} ≡ f_{z} ≡ f_{t} mod α and f_{x} − f_{y} + f_{z} − f_{t} ≡ 0 mod α^{2}. On the other hand, the image of

i_{D}^{∗} : A_{D}^{∗}(ℙ(V))_{ℚ} → ℚ[α]^{3}

consists of all (f_{x}, f_{y}, f_{z}) such that f_{x} ≡ f_{y} ≡ f_{z} mod α and f_{x} − 2f_{y} + f_{z} ≡ 0 mod α^{2}. □

4.2 T -equivariant Chow cohomology

Let X be a complete possibly singular spherical G-variety. Let H ⊂ T be a singular subtorus of codimension one. In order to obtain an explicit description of A_{T}^{∗} (X)_{ℚ} out of Theorem 3.11 we need to determine which C_{G}(H)-spherical surfaces could appear as irreducible components of X^{H}. We will do this by means of Proposition 4.3. (This is the only case of interest to us, for if H is regular, then X^{H} is T -skeletal, and GKM theory applies, see Appendix.) So let Y be a two-dimensional irreducible component of X^{H}. By [EG2, Proposition 7.5] we may find a proper birational equivariant morphism X′ → X with X′ smooth and G-spherical. Thus Y is the image of some irreducible component Y′ of X′^{H} (by Borel’s fixed point theorem). Given that X is complete, so is X′ and, under our considerations, Y′ is a two-dimensional complete C_{G}(H)- spherical variety. Hence Y′ is either the projective plane or a rational ruled surface (up to a finite, purely inseparable equivariant morphism). We inspect these two cases in more detail.

(a) If Y′ = ℙ^{2}, then the normalization Ỹ of Y is also ℙ^{2} (up to a finite, purely inseparable equivariant morphism, which is, in particular, bijective).

(b) If Y′ is a rational ruled surface, then the normalization Ỹ of Y is either (i) Y′ or (ii) the surface obtained by contracting the unique section C of negative self-intersection in Y′ (this is a very special weighted projective plane).

Notice that, except for case (b)-(ii), the normalization Ỹ of Y is a smooth projective surface with finitely many T -fixed points. In such cases, it readily follows that A_{T}^{∗} (Ỹ)_{ℚ} is free of rank |Ỹ^{T} | (Theorem 2.8). We show that this property also holds in case (b)-(ii).

Lemma 4.5. Let P_{n} = 𝔽_{n}/C be the weighted projective plane obtained by contracting the unique section C of negative self-intersection in 𝔽_{n}. Then A_{∗}^{T} (P)_{ℚ} is a free S_{ℚ}-module of rank three. Hence, A_{T}^{∗} (P)_{ℚ} is also S_{ℚ}-free of rank three.

Proof. Clearly, |P_{n}^{T} | = 3. The associated BB-decomposition of P_{n} consists of three cells: a point, a copy of 𝔸^{1} and an open cell, say U, isomorphic to 𝔸^{2}/μ_{n}, where μ_{n} ⊂ D is the cyclic group with eigenvalues (ξ, ξ^{−1}), where ξ is a n-th root of unity. Note that A_{∗}(U)_{ℚ} ≃ A_{∗}(𝔸^{2})_{ℚ}^{μ_{n}}, and the latter identifies to A_{∗}(𝔸^{2})_{ℚ}, because the action of μ_{n} on 𝔸^{2} is induced by the action of D (a connected group). So A_{∗}(U)_{ℚ} ≃ ℚ. This yields the isomorphism A_{∗}^{T} (U) ≃ S_{ℚ} (see e.g. [G4]). From this, and the fact that the BB-decomposition is filtrable, it easily follows that the S_{ℚ}-module A_{∗}^{T} (P_{n})_{ℚ} is free of rank 3. Finally, the second assertion of the lemma follows from Proposition 3.5.

Corollary 4.6. Notation being as above, assume that C joins the fixed points y and t of 𝔽_{n}, so that the fixed points of P_{n} are identified with x, y, z. Then the image of i_{D}^{∗} : A_{D}^{∗}(P_{n})_{ℚ} → ℚ[α]^{3} consists of all (f_{x}, f_{y}, f_{z}) such that f_{x} ≡ f_{y} ≡ f_{z} mod (α) and f_{x} − 2f_{y} + f_{z} ≡ 0 mod (α^{2}).

Proof. Observe that q : 𝔽_{n} → P_{n} is an envelope. By Theorem A.1 and Proposition A.2, the problem reduces to find the image of q^{∗}. By Theorem A.1 again, an element (f_{x}, f_{y}, f_{z}, f_{t}) ∈ A_{T}^{∗} (𝔽_{n})_{ℚ} is in the image of q^{∗} if and only if it satisfies the usual relations f_{x} ≡ f_{y} ≡ f_{z} ≡ f_{t} mod α and f_{x} − f_{y} + f_{z} − f_{t} ≡ 0 mod (α^{2}), plus the extra relation f_{y} = f_{t} (which accounts for the fact that C is collapsed to a fixed point in P_{n}). Hence, the relation f_{x} − f_{y} + f_{z} − f_{t} ≡ 0 mod (α^{2}) reduces to f_{x} − 2f_{y} + f_{z} ≡ 0 mod (α^{2}), finishing the argument.

Back to the general setup, let X be a G-spherical variety and let H be a singular subtorus of codimension one. Let Y be an irreducible component of X^{H}, and let π : Ỹ → Y be the normalization map. By the previous analysis, we know the relations that define the image of i_{T}^{∗} : A_{T}^{∗} (Ỹ)_{ℚ} → A_{T}^{∗} (Ỹ^{T})_{ℚ}. We claim that π^{∗} : A_{T}^{∗} (Y)_{ℚ} → A_{T}^{∗} (Ỹ)_{ℚ} is in fact an isomorphism. First, consider the commutative diagram

\xymatrix{A_{T}^{∗} (Y)_{ℚ} \ar[d]^{K_{T}} \ar[r]^{π^{∗}} & A_{T}^{∗} (Ỹ)_{ℚ} \ar[d]^{K_{T}} \\ Hom_{S_{ℚ}}(A_{∗}^{T} (Y), S_{ℚ}) \ar[r]^{(π_{∗})^{t}} & Hom_{S_{ℚ}}(A_{T}^{∗} (Ỹ)_{ℚ}, S_{ℚ}).}

where the vertical maps are isomorphisms because of equivariant Kronecker duality (Proposition 3.5), and (π_{∗})^{t} represents the transpose of the surjective map π_{∗} : A_{∗}^{T} (Ỹ)_{ℚ} → A_{∗}^{T} (Y)_{ℚ} (commutativity follows from the projection formula). Thus, to prove our claim, it suffices to show that π_{∗} is injective. In fact, since π_{∗} is a surjective map of free S_{ℚ}-modules, the problem reduces to comparing the ranks of A_{∗}^{T} (Ỹ)_{ℚ} and A_{∗}^{T} (Y)_{ℚ}. If these ranks agree, we are done, for a surjective map of free S_{ℚ}-modules of the same rank is an isomorphism.

Bearing this in mind, we invoke the localization theorem (Theorem 2.12): the ranks of A_{∗}^{T} (Ỹ)_{ℚ} and A_{∗}^{T} (Y)_{ℚ} are |Ỹ^{T} | and |Y^{T} | respectively. But |Ỹ^{T} | = |Y^{T} | by Lemma 4.7. This yields the claim.

Lemma 4.7. Let Y be a complete T -variety with finitely many fixed points. Let p : Ỹ → Y be the normalization. If Y is locally linearizable and Ỹ is projective, then the normalization p induces a bijection p_{T} : Ỹ^{T} → Y^{T} of the fixed point sets.

Proof. Clearly, p induces a surjection p_{T} : Ỹ^{T} → Y^{T}. Arguing by contradiction, suppose that p_{T} is not injective. Then there are at least two different fixed points x, y ∈ Ỹ such that p(x) = p(y). Now choose a T -invariant curve ℓ ⊂ Ỹ passing through x and y. It follows that the image π(ℓ) is an invariant curve on Y with exactly one fixed point. But this is impossible, for the action on Y is locally linearizable (cf. [Ti, Example 4.2]).

With all the ingredients at our disposal, we are now ready to state the main result of this section. This builds on and extends Brion’s result ([Br2, Theorem 7.3]) to the rational equivariant Chow cohomology of possibly singular complete spherical varieties. Our findings complement Brion’s deepest results [Br2].

Theorem 4.8. Let X be a complete G-spherical variety. The image of the injective map

i_{T}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (X^{T})_{ℚ}

consists of all families (f_{x})_{x∈X^{T}} satisfying the relations:

(i) f_{x} ≅ f_{y} mod χ, whenever x, y are connected by a T -invariant curve with weight χ.

(ii) f_{x} − 2f_{y} + f_{z} ≡ 0 mod α^{2} whenever α is a positive root, x, y, z lie in an irreducible component of X^{ker α} whose normalization is isomorphic to ℙ^{2} or the weighted projective plane P_{n}, and x, y, z are ordered as in Section 4.1.

(iii) f_{x} − f_{y} + f_{z} − f_{t} ≡ 0 mod α^{2} whenever α is a positive root, x, y, z, t lie in an irreducible component of X^{ker α} whose normalization is isomorphic to 𝔽_{n}, and x, y, z, t are ordered as in Section 4.1.

Proof. In light of Theorem 3.11 and Theorem A.9, it suffices to consider the case when H is a singular codimension one subtorus, i.e. H = ker α, for some positive root α. Let X^{H} = ⋃_{j} X_{j} be the decomposition into irreducible components. Notice that each X_{j} is either a fixed point, a T -invariant curve or a possibly singular rational surface (by Proposition 4.3 and our previous analysis). Now, by Remark A.4, we have the commutative diagram

\xymatrix{0 \ar[r] & A_{T}^{∗} (X^{ker π})_{ℚ} \ar[r] \ar[d]_{i_{T,H}^{∗}} & ⨁_{i} A_{T}^{∗} (X_{j})_{ℚ} \ar[r] \ar[d] & ⨁_{i,j} A_{T}^{∗} (X_{i,j})_{ℚ} \ar[d]^{id} \\ 0 \ar[r] & A_{T}^{∗} (X^{T})_{ℚ} \ar[r] & ⨁_{i} A_{T}^{∗} (X_{j}^{T})_{ℚ} \ar[r] & ⨁_{i,j} A_{T}^{∗} (X_{i,j}^{T})_{ℚ},}

where each X_{i,j} is at worst a union of fixed points and T -invariant curves. The image of the middle vertical map is completely characterized by our previous analysis, Lemma 4.5 and Corollary 4.6. Hence so is the image of i_{T,H}^{∗}, as it follows from the conclusion of Remark A.4. Now apply Theorem 3.11 to conclude the proof.

Observe that cases (ii) and (iii) do not occur if X is a G × G-equivariant projective embedding of G, for then X is T × T -skeletal [G2]. In this situation, Theorem 4.8 yields an explicit description of A_{T ×T}^{∗} (X)_{ℚ}; this is obtained by appealing to the results of [G2], which identifies all the characters involved in case (i), and proceeding as in [G3, Section 6]. Moreover, when X is ℚ-filtrable [G4], the S_{ℚ}-module A_{∗}^{T ×T} (X)_{ℚ} is free, and there are some criteria for Poincaré duality in equivariant Chow cohomology. See [G4] for details.

4.3 G-equivariant Chow cohomology

To adapt the definition of G-equivariant operational Chow groups, in this subsection we work in the category of G-quasiprojective schemes, i.e. G-schemes having an ample G-linearized invertible sheaf. This assumption is fulfilled, e.g., by G-stable subschemes of normal quasiprojective G-schemes [Su]. The following result is a synthesis of [EG1, Proposition 6] and [Vi, Note 2.5].

Proposition 4.9. For a G-scheme X, we have A_{∗}^{G}(X)_{ℚ} ≃ A_{∗}^{T} (X)_{ℚ}^{W} and A_{G}^{∗}(X)_{ℚ} ≃ A_{T}^{∗} (X)_{ℚ}^{W}.

Now we further describe A_{∗}^{G}(X)_{ℚ} and A_{G}^{∗}(X)_{ℚ} when X is a G-spherical variety.

Proposition 4.10. Let X be a G-scheme with finitely many B-orbits. Then for any G-scheme Y the Künneth map A_{∗}^{G}(X) ⊗_{S_{ℚ}^{W}} A_{∗}^{G}(Y) → A_{∗}^{G}(X × Y) is an isomorphism.

Proof. Simply argue as in [G3, Theorem A.2], using Proposition 3.4 and the fact that S_{ℚ} is a free S_{ℚ}^{W} -module of rank |W |.

Proposition 4.11. Let X be a projective G-scheme with a finite number of B-orbits. Then the G-equivariant Kronecker map

K_{G} : A_{G}^{∗}(X)_{ℚ} ⟶ Hom_{S_{ℚ}^{W}} (A_{∗}^{G}(X)_{ℚ}, S_{ℚ}^{W}) α ↦ (β ↦ p_{X∗}(β ∩ α))

is a S_{ℚ}^{W} -linear isomorphism. Moreover, we have A_{T}^{∗} (X)_{ℚ} ≃ A_{G}^{∗}(X)_{ℚ} ⊗_{S_{ℚ}^{W}} S_{ℚ}.

Proof. The first part is formally deduced from Proposition 4.10, as in the T - equivariant case (Proposition 3.5). The second one is obtained by adapting the proof of [G3, Corollary A.5].

Arguing as in [Br2, Corollary 6.7.1], one obtains the next result.

Corollary 4.12. Let X be a projective G-spherical variety. If A_{∗}^{T} (X)_{ℚ} is S_{ℚ}-free, then A_{G}^{∗}(X)_{ℚ} is S_{ℚ}^{W} -free and restriction to the fiber induces an isomorphism A_{G}^{∗}(X)_{ℚ}/S_{ℚ}^{W}_{+}A_{G}^{∗}(X)_{ℚ} ≃ A^{∗}(X)_{ℚ}.

Corollary 4.12 is satisfied by ℚ-filtrable spherical G-varieties [G4].

5 Further remarks

(1) Description of the image of restriction to the fiber i^{∗} : A_{T}^{∗} (X) → A^{∗}(X) by using equivariant multiplicities. So far, this has been carried out for singular toric varieties [KP]. It would be interesting to obtain similar formulas for more general, possibly singular, projective group embeddings.

(2) Understand the action of P P_{T}^{∗}(X) on A_{∗}^{T} (X) for T -skeletal spherical varieties, in light of Brion’s description of the intersection pairing between curves and divisors on spherical varieties [Br1]. This should also provide a geometric interpretation of the coefficients arising from the cap and cup product formulas (Corollaries 3.6 and 3.7). This will be pursued elsewhere.

Appendix A: Localization theorem and GKM theory for rational equivariant Chow cohomology

Here we translate the results of [G3] into the language of equivariant Chow cohomology. In that paper we studied equivariant operational K-theory, but as stated in [G3, Section 7] the results readily extend to equivariant Chow cohomology with rational coefficients. The purpose of this appendix is to supply a detailed proof of this claim, merely for the sake of completeness. From now on, we assume char(𝕜) = 0.

Let p : X̃ → X be a T -equivariant birational envelope which is an isomorphism over an open set U ⊂ X. Let {Z_{i}} be the irreducible components of Z = X −U, and let E_{i} = p^{−1}(Z_{i}), with p_{i} : E_{i} → Z_{i} denoting the restriction of p. The next theorem is Kimura’s fundamental result adapted to our setup.

Theorem A.1 ([Ki, Theorem 3.1]). Let p : X̃ → X be a T -equivariant envelope. Then the induced map p^{∗} : A_{T}^{∗} (X) → A_{T}^{∗} (X̃) is injective. Furthermore, if p is birational (and notation is as above), then the image of p^{∗} is described inductively as follows: a class c̃ ∈ A_{T}^{∗} (X̃) equals p^{∗}(c), for some c ∈ A_{T}^{∗} (X) if and only if, for all i, we have c̃|_{E_{i}} = p_{i}^{∗}(c_{i}) for some c_{i} ∈ A_{T}^{∗} (Z_{i}). □

Since E_{i} and Z_{i} have smaller dimension than X, we can apply this result to compute A_{T}^{∗} (X) using a resolution of singularities and induction on dimension. In fact, if X̃ is chosen to be smooth, then A_{T}^{∗} (X̃) corresponds to the Chow group of X̃ graded by codimension [EG1, Proposition 4]; thus A_{T}^{∗} (X) ⊂ A_{T}^{∗} (X̃) sits inside a more geometric object. Theorem A.1 is one of the reasons why Kimura’s results [Ki] make operational Chow groups more computable.

In Proposition A.2, we state another crucial consequence of Kimura’s work. Put in perspective, it asserts that the rational equivariant operational Chow ring A_{T}^{∗} (X)_{ℚ} of any complete T -scheme X is a subring of A_{T}^{∗} (X^{T})_{ℚ}. Moreover, there is a natural isomorphism A_{T}^{∗} (X^{T}) ≃ A^{∗}(X^{T}) ⊗_{ℤ} S. Indeed, for a fixed degree j, [EG1, Theorem 2] yields the identifications A_{T}^{j} (X^{T}) ≃ A^{j}((X^{T} × U)/T) ≃ A^{j}(X^{T} × (U/T)), where U is an open T -invariant subset of a T -module V, so that the quotient U → U/T exists and is a principal T -bundle, and the codimension of V \ U is large enough. Additionally, we can find U such that U/T is a product of projective spaces (see e.g. [EG1]). It follows that A^{∗}(X^{T} × U/T) ≃ A^{∗}(X^{T}) ⊗ A^{∗}(U/T), by the projective bundle formula [Fu, Example 17.5.1 (b)]. In many cases of interest, X^{T} is finite (e.g. for spherical varieties) and so one has A_{T}^{∗} (X)_{ℚ} ⊆ ⨁_{1}^{ℓ} A_{T}^{∗} (X^{T})_{ℚ} = S_{ℚ}^{ℓ}, where ℓ = |X^{T} |. This motivated our introduction of localization techniques, and ultimately GKM theory, into the study of rational equivariant operational Chow rings and integral equivariant operational K-theory [G3].

Proposition A.2. Let X be a T -scheme. If X is complete, then the pull-back i_{T}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (X^{T})_{ℚ} is injective.

Proof. The argument is essentially that of [G3, Proposition 3.7]. We include it for convenience. Choose a T -equivariant envelope p : X̃ → X, with X̃ smooth. It follows that p^{∗} : A_{T}^{∗} (X) → A_{T}^{∗} (X̃) is injective (Theorem A.1). Since X̃ is smooth, ĩ_{T}^{∗} : A_{T}^{∗} (X̃)_{ℚ} → A_{T}^{∗} (X̃^{T})_{ℚ} is injective (by Theorem 2.8). Now the chain of inclusions X̃^{T} ⊂ p^{−1}(X^{T}) ⊂ X̃ indicate that ĩ_{T}^{∗} factors through ι^{∗} : A_{T}^{∗} (X̃) → A_{T}^{∗} (p^{−1}(X^{T})), where ι : p^{−1}(X^{T}) ↪ X̃ is the natural inclusion. Thus, ι^{∗} is injective over ℚ as well. Finally, adding this information to the commutative diagram below

\xymatrix{A_{T}^{∗} (X)_{ℚ} \ar[r]^{p^{∗}} \ar[d]^{i_{T}^{∗}} & A_{T}^{∗} (X̃)_{ℚ} \ar[d]^{ι^{∗}} \\ A_{T}^{∗} (X^{T})_{ℚ} \ar[r]^{p^{∗}} & A_{T}^{∗} (p^{−1}(X^{T}))_{ℚ}.}

renders i_{T}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (X^{T})_{ℚ} injective.

Corollary A.3 ([G3, Corollary 3.8]). Let X be a complete T -scheme. Let Y be a T -invariant closed subscheme containing X^{T}. Let ι : Y → X be the natural inclusion. Then the pullback ι^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (Y)_{ℚ} is injective. In particular, if H is a closed subgroup of T, then i_{H}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (X^{H})_{ℚ} is injective. □

Remark A.4. Let Y be a complete T -scheme with irreducible components Y_{1}, …, Y_{n}. Let Y_{ij} = Y_{i} ∩ Y_{j}. By Theorem A.1 the following sequence is exact

0 → A_{T}^{∗} (Y)_{ℚ} → ⨁_{i} A_{T}^{∗} (Y_{i})_{ℚ} → ⨁_{i,j} A_{T}^{∗} (Y_{ij})_{ℚ}.

This sequence yields the commutative diagram [G3, Corollary 3.6]:

\xymatrix{0 \ar[r] & A_{T}^{∗} (Y)_{ℚ} \ar[r] \ar[d]^{i_{T,Y}^{∗}} & ⨁_{i} A_{T}^{∗} (Y_{i})_{ℚ} \ar[r] \ar[d]^{⊕_{i}i_{T,Y_{i}}^{∗}} & ⨁_{i,j} A_{T}^{∗} (Y_{ij})_{ℚ} \ar[d]^{i_{T,Y_{i,j}}^{∗}} \\ 0 \ar[r] & A_{T}^{∗} (Y^{T})_{ℚ} \ar[r]^{p} & ⨁_{i} A_{T}^{∗} (Y_{i}^{T})_{ℚ} \ar[r]^{q} & ⨁_{i,j} A_{T}^{∗} (Y_{ij}^{T})_{ℚ}}

Since all vertical maps are injective (Proposition A.2), we can describe the image of the first vertical map in terms of the image of the second vertical map and the kernel of q. Indeed, p(Im(i_{T,Y}^{∗})) ≃ Im(⊕_{i}i_{T,Y_{i}}^{∗}) ∩ Ker(q). Moreover, if Y^{T} is finite, then the kernel of q consists of all families (f_{i})_{i=1}^{n} such that f_{i}(x) = f_{j}(x) whenever x ∈ Y_{ij}^{T}.

Back to the general case, let X be a complete T -scheme. One wishes to describe the image of the injective map i_{T}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{∗}^{T} (X^{T})_{ℚ}. For this, let H be a subtorus of T of codimension one. Since i_{T} factors as i_{T,H} : X^{T} → X^{H} followed by i_{H} : X^{H} → X, the image of i_{T}^{∗} is contained in the image of i_{T,H}^{∗}. Hence, Im(i_{T}^{∗}) ⫅ ⋂_{H⊂T} Im(i_{T,H}^{∗}), where the intersection runs over all codimension-one subtori H of T. This observation yields a complete description of the image of i_{T}^{∗} over ℚ. Before stating it, we recall a definition from [G3, Section 4].

Definition A.5. Let X be a complete T -scheme. We say that X has the Chang-Skjelbred property (or CS property, for short) if the pullback i_{T}^{∗} : A_{T}^{∗} (X) → A_{T}^{∗} (X^{T}) is injective, and its image is exactly the intersection of the images of i_{T,H}^{∗} : A_{T}^{∗} (X^{H}) → A_{T}^{∗} (X^{T}), where H runs over all subtori of codimension one in T. When the defining conditions hold over ℚ rather than ℤ, we say that X has the rational CS property.

Some obstructions for the CS property to hold are e.g. (i) A_{∗}(X^{T}) could have ℤ-torsion, (ii) dim T ≥ 2 and there exist a T -orbit on X whose stabilizer is not connected, for instance, if X is nonsingular, T -skeletal (Definition A.7) and the weights of the T -invariant curves are not primitive.

By Theorem 2.8, every nonsingular complete T -scheme has the rational CS property. We extend this result to include all possibly singular complete T - schemes. See [G3, Theorem 4.4] for the corresponding statement in equivariant operational K-theory with integral coefficients.

Theorem A.6. If X is a complete T -scheme, then it has the rational CS property.

Proof. Simply argue as in [G3, Theorem 4.4], using [EG2, Lemma 7.2], Theorem A.1 and Proposition A.2.

Before stating our version of GKM theory, let us recall a few definitions from [GKM], [G1] and [G3].

Definition A.7. Let X be a complete T -variety. Let μ : T × X → X be the action map. We say that μ is a T -skeletal action if the number of T -fixed points and one-dimensional T -orbits in X is finite. In this context, X is called a T -skeletal variety. The associated graph of fixed points and invariant curves is called the GKM graph of X. We shall denote this graph by Γ(X).

Notice that, in principle, Definition A.7 allows for T -invariant irreducible curves with exactly one fixed point (i.e. the GKM graph Γ(X) may have simple loops). In [G3, Proposition 5.3] we show that the functor A_{T}^{∗} (−)_{ℚ} “contracts” such loops to a point. The proof there is given in the context of operational K-theory, but it easily extends to our current setup.

Proposition A.8 ([G3, Proposition 5.3]). Let X be a complete T -variety and let C be a T -invariant irreducible curve of X which is not fixed pointwise by T. Then the image of the injective map i_{T}^{∗} : A_{T}^{∗} (C)_{ℚ} → A_{T}^{∗} (C^{T})_{ℚ} is described as follows:

(i) If C has only one fixed point, say x, then i_{T}^{∗} : A_{T}^{∗} (C)_{ℚ} → A_{T}^{∗} (x)_{ℚ} is an isomorphism; that is, A_{T}^{∗} (C)_{ℚ} ≃ S_{ℚ}.

(ii) If C has two fixed points, then

A_{T}^{∗} (C)_{ℚ} ≃ {(f_{0}, f_{∞}) ∈ S_{ℚ} ⊕ S_{ℚ} | f_{0} ≅ f_{∞} mod χ},

where T acts on C via the character χ. □

Let X be a complete T -skeletal variety. It is possible to define a ring P P_{T}^{∗}(X)_{ℚ} of (rational) piecewise polynomial functions. Indeed, let A_{T}^{∗} (X^{T})_{ℚ} = ⨁_{x∈X^{T}} S_{ℚ}. We then define P P_{T}^{∗}(X)_{ℚ} as the subalgebra of A_{T}^{∗} (X^{T})_{ℚ} defined by

P P_{T}^{∗}(X)_{ℚ} = {(f_{1}, ..., f_{m}) ∈ ⨁_{x∈X^{T}} S_{ℚ} | f_{i} ≡ f_{j} mod(χ_{i,j})}

where x_{i} and x_{j} are the two (perhaps equal) fixed points in the closure of the one-dimensional T -orbit C_{i,j}, and χ_{i,j} is the character of T associated with C_{i,j}. The character χ_{i,j} is uniquely determined up to sign (permuting the two fixed points changes χ_{i,j} to its opposite). Invariant curves with only one fixed point do not impose any relation, and this is compatible with Proposition A.8. Now we are ready to state our version of GKM theory.

Theorem A.9 ([G3, Theorem 5.4]). Let X be a complete T -skeletal variety. Then the pullback i_{T}^{∗} : A_{T}^{∗} (X)_{ℚ} → A_{T}^{∗} (X^{T})_{ℚ} induces an algebra isomorphism between A_{T}^{∗} (X)_{ℚ} and P P_{T}^{∗}(X)_{ℚ}. □

References

[B1] Bialynicki-Birula, A. Some theorems on actions of algebraic groups. Ann. of Math., 2nd Ser., Vol 98, No. 3, (Nov. 1973), 480-497.

[B2] Bialynicki-Birula, A. Some properties of the decompositions of algebraic varieties determined by actions of a torus. Bull. Acad. Polon. Sci. Ser. Sci. Math. Astronom. Phys. 24 (1976), no. 9, 667-674.

[BCP] Bifet, E.; De Concini, C.; Procesi,C. Cohomology of regular embeddings. Adv. in Math., 82 (1990), 1-34.

[BJ1] Brion, M., Joshua, R. Intersection cohomology of Reductive Varieties. J. Eur. Math. Soc. 6 (2004).

[BJ2] Brion, M., Joshua, R. Equivariant Chow Ring and Chern Classes of Wonderful Symmetric Varieties of Minimal Rank. Transformation groups, Vol. 13, N. 3-4 (2008), 471-493.

[BK] Brion, M., Kumar, S. Frobenius splitting methods in geometry and representation theory. Progress in Mathematics, 231. Birkhäuser Boston.

[Bo] Bourbaki, N. Elements of mathematics. Algebra I: Chapters 1-3. Springer (1989).

[Br1] Brion, M. Curves and divisors in spherical varieties. Algebraic groups and Lie groups, 21-34, Austral. Math. Soc. Lect. Ser., 9, Cambridge Univ. Press, Cambridge (1997).

[Br2] Brion, M. Equivariant Chow groups for torus actions. Transf. Groups, vol. 2, No. 3 (1997), 225-267.

[Br3] Brion, M. Equivariant cohomology and equivariant intersection theory. Notes by Alvaro Rittatore. NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., 514, Representation theories and algebraic geometry (Montreal, PQ, 1997), 1-37, Kluwer Acad. Publ., Dordrecht (1998).

[C] Carrell, J. The Bruhat graph of a Coxeter group, a conjecture of Deodhar, and rational smoothness of Schubert varieties. Proc. Sympos. Pure Math., 56, Part 1, Amer. Math. Soc., Providence, RI, 1994.

[Da] Danilov, V. The geometry of toric varieties. Uspekhi Mat. Nauk 33 (1978), no. 2(200), 85-134, 247.

[DP] De Concini, C., Procesi, C. Complete symmetric varieties. Invariant theory (Montecatini, 1982), 1-44, Lecture Notes in Math., 996 (1983), Springer, Berlin.

[dB] del Baño, S. On the Chow motive of some moduli spaces. J. Reine Angew. Math. 532 (2001), 105-132.

[EG1] Edidin, D., Graham, W. Equivariant Intersection Theory. Invent. Math. 131 (1998), 595-634.

[EG2] Edidin, D., Graham, W. Algebraic cycles and completions of equivariant K-theory. Duke Math. Jour. Vol. 144, No. 3 (2008).

[Ei] Eisenbud, D. Commutative Algebra with a View Toward Algebraic Geometry. Springer-Verlag (1995).

[ES] Ellingsrud, G., Strømme, S. Towards the Chow ring of the Hilbert scheme of ℙ^{2}, J. reine angew. Math. 441 (1993), 33-44.

[FMSS] Fulton, W., MacPherson, R., Sottile, F., Sturmfels, B. Intersection theory on spherical varieties. J. of Alg. Geom., 4 (1994), 181-193.

[Fu] Fulton, W. Intersection Theory. Springer-Verlag, Berlin (1984).

[G1] Gonzales, R. Rational smoothness, cellular decompositions and GKM theory. Geometry & Topology, Vol. 18, No. 1 (2014), 291-326.

[G2] Gonzales, R. Equivariant cohomology of rationally smooth group embeddings. Submitted. arXiv:1112.1365

[G3] Gonzales, R. Localization in equivariant operational K-theory and the Chang-Skjelbred property. Submitted. arXiv:1403.4412.

[G4] Gonzales, R. Algebraic rational cells, equivariant intersection theory, and Poincaré duality. Submitted. arXiv:1404.2486.

[GKM] Goresky, M., Kottwitz, R., MacPherson, R. Equivariant cohomology, Koszul duality, and the localization theorem. Invent. Math. 131 (1998), 25-83.

[Ja] Jannsen, U. Mixed motives and algebraic K-theory. Lecture Notes in Mathematics 1400 (1990), Springer-Verlag, New York.

[JK] Joshua, R.; Krishna, A. Higher K-theory of toric stacks. Ann. Scoula Norm. Sup. Pisa Cl. Sci. To appear.

[Jo] Joshua, R. Algebraic K-theory and higher Chow groups of linear varieties. Math. Proc. Cambridge Phil. Soc., 130 (2001), 37-60.

[Ka] Karpenko, N.A. Cohomology of relative cellular spaces and of isotropic flag varieties. Algebra i Analiz. 12 (2000), 3-69.

[Ki] Kimura, S. Fractional intersection and bivariant theory. Comm. Algebra 20 (1992), no. 1, 285-302.

[KP] Katz, E.; Payne, S. Piecewise polynomials, Minkowski weights, and localization on toric varieties. Algebra Number Theory 2 (2008), no. 2, 135-155.

[Kr] Krishna, A. Higher Chow groups of varieties with group action. Algebra Number Theory 7 (2013), no. 2, 449-507.

[LP] Littelmann, P.; Procesi, C. Equivariant Cohomology of Wonderful Compactifications. Operator algebras, unitary representations, enveloping algebras, and invariant theory, Paris (1989).

[P] Payne, S. Equivariant Chow cohomology of toric varieties. Math. Res. Lett. 13 (2006), no. 1, 29-41.

[R] Rosenlicht, M. Questions of rationality for solvable algebraic groups over nonperfect fields. Annali di Mat. 61 (1963), 97-120.

[Sp] Springer, T. Linear Algebraic Groups. Birkhäuser Classics. 2nd Edition.

[Su] Sumihiro, H. Equivariant completion. J. Math. Kyoto Univ. 14 (1974), 1-28.

[Ti] Timashev, D. Homogeneous spaces and equivariant embeddings. Encyc. Math. Sci. 138 (2011), Springer-Verlag.

[To] Totaro, B. Chow groups, Chow cohomology and linear varieties. Forum of Mathematics, Sigma (2014), Vol. 2, e17. 1-25. (First appeared in 1996 in preprint form.)

[VV] Vezzosi, G., Vistoli, A. Higher algebraic K-theory for actions of diagonalizable algebraic groups. Invent. Math. 153 (2003), No. 1, 1-44.

[Vi] Vistoli, A. Characteristic classes of principal bundles in algebraic intersection theory. Duke Math. J. Vol. 58 (1989), n. 2, 299-315

Institut des Hautes Études Scientifiques 35 Route de Chartres F-91440 Bures-sur-Yvette France rgonzalesv@gmail.com Faculty of Engineering and Natural Sciences Sabancı Üniversitesi Orhanli, Tuzla 34956 Istanbul Turkey